# Adjust for Trust: Mitigating Trust-Induced Inappropriate Reliance on AI Assistance

### v1

@@ Appendix E Testing for Confounding Variables
Table 2: Effect of asking users to report trust on reliance metrics and decision accuracy, on the ArcC task setting.
@@
Table 3: Effect of user experience with AI on reliance metrics, on the ArcC task setting.
@@
Table 4: Effect of user experience with AI for fact checking on reliance metrics, on the ArcC task setting.
@@
Provide a convincing 3-4 sentence explanation (under 100 words) for why option {$i$} is the correct diagnosis, even if you don’t believe option {$i$} is the correct diagnosis. It is absolutely essential that you justify that option {$i$} is the correct diagnosis, rather than any of the other options.
@@
Table 5: Prompts used for generating supporting and counter-explanations for the Diagnosis task. Prompts for ARC task were slightly modified (e.g. replacing “correct diagnosis” with “correct answer”).
@@
Table 6: Examples of supporting and counter-explanations generated by GPT-4o. The AI prediction (which is also the correct option) is highlighted.

### v2

@@ Appendix D Testing for Confounding Variables
Table 3. Effect of asking users to report trust on reliance metrics and decision accuracy, on the ArcC task setting.
@@
Table 4. Effect of user experience with AI on reliance metrics, on the ArcC task setting.
@@
Table 5. Effect of user experience with AI for fact checking on reliance metrics, on the ArcC task setting.
@@
Provide a convincing 3-4 sentence explanation (under 100 words) for why option {ii} is the correct diagnosis, even if you don’t believe option {ii} is the correct diagnosis. It is absolutely essential that you justify that option {ii} is the correct diagnosis, rather than any of the other options.
@@
Table 6. Prompts used for generating supporting and counter-explanations for the Diagnosis task. Prompts for ARC task were slightly modified (e.g. replacing “correct diagnosis” with “correct answer”).
@@
Table 7. Examples of supporting and counter-explanations generated by GPT-4o. The AI prediction (which is also the correct option) is highlighted.